### v2

# MAPO: Momentum-Aided Gradient Descent Prompt Optimization

Abstract
Momentum-Aided Prompt Optimization (MAPO) enhances the efficiency and efficacy of prompt optimization for Large Language Models (LLMs). Building on ProTeGi (Pryzant et al., 2023), MAPO uses positive natural language "gradients" and a momentum-based extension to refine prompts effectively. By tracking gradient history, MAPO avoids local minima and oscillations. It also utilizes beam search and an Upper Confidence Bound (UCB) algorithm for balanced candidate expansion and selection. Benchmark testing shows that MAPO achieves faster convergence time with fewer API calls and higher F1 scores than ProTeGi, proving it as a robust and scalable solution for automated prompt engineering in LLMs.
MAPO: Momentum-Aided Gradient Descent Prompt Optimization
Anthony Cui           Pranav Nandyalam           Ethan Cheung           Kevin Zhu Algoverse AI Research acui@overlake.org, pranavrajnandyalam@gmail.com, kevin@algoverse.us
1 Introduction
Large Language Models (LLMs) have gained significant attention since the release of ChatGPT OpenAI (2022), leading to the development of new prompting techniques that have greatly improved LLM performance (Schulhoff et al., 2024). However, prompts can still be unclear, biased, or incomplete, limiting LLM capabilities (Sahoo et al., 2024). Prompt engineering has become critical, but current methods often require manual adjustments, making them time-consuming, error-prone, and constrained by human limitations (Lin et al., 2024). This highlights an increasing need for an automated system to improve prompt quality.
Recent work has explored implementing traditional machine learning algorithms into a natural language format, with one of the first being ProTeGi’s “Automatic Prompt Optimization with ‘Gradient Descent’ and Beam Search” (Pryzant et al., 2023). While ProTeGi introduced an innovative framework, it has limitations such as excessive API calls, resource consumption, and underutilizing the strengths of prompts.
Figure 1: High-Level Overview of MAPO
We introduce Momentum-Aided Prompt Optimization (MAPO), a method that extends ProTeGi by using positive natural language “gradients” with momentum to automate prompt refinement. Gradients are generated from correct examples in a minibatch, guiding the LLM to refine prompts in a consistent semantic direction. Beam search expands the candidate pool, and the best-arm identification algorithm, using UCB bandits, selects the top prompts for further evaluation.
MAPO improves on ProTeGi by speeding up convergence and reducing resource use through momentum-based adjustments. By tracking gradient history, MAPO avoids local minima and oscillations. In a case study, MAPO demonstrates faster runtime, higher F1 scores, and fewer API calls, providing a scalable solution for automated prompt engineering in LLMs.
2 Related Works
Prompt Engineering. In this work, we draw upon existing prompt engineering techniques and focus on incorporating optimization algorithms into our framework to enhance the effectiveness of prompt optimization. There is an increasingly diverse set of general frameworks that previous works have focused on: LLM optimization ((Pryzant et al., 2023); (Zelikman et al., 2023); (Fernando et al., 2023); (Zhou et al., 2022); (Yang et al., 2023)), reinforcement learning ((Ma et al., 2023); (Zhang et al., 2022); (Deng et al., 2022)), and in-context learning (Shum et al., 2023). However, these approaches are generally not feasible when there is no architectural information introduced and only an API is provided to the LLM. More specifically, we base most of our work on improving automatic prompt engineering techniques with LLM gradient-based methods ((Shin et al., 2020) (Pryzant et al., 2023)). Though, many of the current methods involving some sort of iterative refinement technique, such as APE (Zhou et al., 2022) and ProTeGi (Pryzant et al., 2023) all face similar struggles with cumulative costs of running their programs.
3 Methods
3.1 Momentum-Aided Prompt Optimization
First, the current prompt $\mathbf{p}$ is evaluated based on a minibatch of training data to obtain randomly sampled strings $\mathbf{s}$, which contain correct LLM predictions generated relative to the correct labels. We then provide the LLM with a static prompt $\tau$ to generate numerous positive “gradients” $\nabla\mathbf{p}$ in natural language, praising the current prompt $\mathbf{p}$ using the sampled strings $\mathbf{s}$. Our “gradients” $\nabla\mathbf{p}$ are the natural language outputs of the LLM’s continuation of static prompt $\tau$. In traditional machine learning, gradient descent uses numerical gradients, representing a vector in parameter space where a model can improve or worsen; by contrast, our textual gradients $\nabla\mathbf{p}$ represent directions in semantic space (Pryzant et al., 2023). We use another static prompt $\alpha$ to apply these gradients to the initial prompt $\mathbf{p}$, allowing us to move along the same semantic direction as the positive textual gradients, refining and improving the initial prompt.
3.2 Expansion and Selection
Our method employs beam search to explore the space of prompt variations generated during optimization. In each round, new candidate prompts $\mathbf{p}_{k}$ are created from the top $k$ best performing prompts from the previous round by iterating through the prompt optimization process described in Section 3.1. After each round, less promising candidates are pruned, and only the top $k$ prompts are retained for further gradient-based improvements, evaluated by a scoring function that assesses how well they meet our predefined objectives, such as F1 score.
We also record the gradients from static prompt $\alpha$, used to generate the top $k$ candidates in each round, to incorporate our novel momentum extension into the natural language gradient descent. Drawing on the physics intuition of momentum, traditional gradient descent uses this extension to improve stability and convergence, helping the model avoid oscillations and escape local minima, thereby reaching global minima more efficiently. Analogously, our method maintains a history of past gradients, guiding the movement of the initial prompt $\mathbf{p}$ in each beam search round through semantic space, helping it converge on the optimal prompt rather than just an incremental improvement. A single positive gradient is randomly sampled from a pool of all the gradients used to generate the top $k$ prompts in each beam search round, representing the positive gradient history. This gradient history is then incorporated into our static prompts $\tau$ and $\alpha$ as textual momentum, guiding the LLM to generate new gradients $\nabla\mathbf{p}$ and new prompt candidates $\mathbf{p}_{c}$, allowing the initial prompt $\mathbf{p}$ to “roll down the hill” faster, i.e., achieve a much faster rate of convergence during prompt optimization.
In our implementation, we utilize the same Upper Confidence Bound (UCB) Bandits Selection algorithm as ProTeGi (Pryzant et al., 2023), which is employed each beam search expansion to evaluate the candidate prompts. UCB has proven to be the strongest best arm identification algorithm for maximizing test metrics such as F1 score, outperforming other algorithms like UCB-E, Successive Rejects, and Successive Halving (Pryzant et al., 2023). UCB effectively balances exploration—searching for better options—and exploitation—selecting the best-performing candidates.
Figure 2: Test performance (F1) versus Time for Liar and Ethos datasets.
Figure 3: Test performance (F1) versus Total API Calls for Liar and Ethos datasets.
Figure 4: Test performance (F1) versus Beam Search Round for Liar and Ethos datasets.
4 Experiments
4.1 Setup
Our experimental setup closely follows that of ProTeGi, allowing for a direct comparison between our extension method and their baseline. We use 200 randomly sampled data points as the test set, retaining most hyperparameters from ProTeGi’s configuration, including a temperature of 0, a minibatch size of 64, and 6 rounds of beam search with a selection size of 4. In each iteration (as outlined in Section 2.1), we generate 2 positive gradients using 3 randomly sampled correct examples from the minibatch, resulting in fewer gradients than ProTeGi’s setup, which uses 4 negative gradients. This trade-off improves runtime efficiency while handling the added complexity of gradient history. The primary evaluation metric is the F1 score, and results reflect the highest score among the top k beam search candidates, averaged over three trial runs to account for variability. Unless otherwise stated, all experiments use the October 2024 release of GPT-3.5-turbo.
4.1.1 Baseline
ProTeGi. Developed by Pryzant et al. (2023), ProTeGi employs natural language gradient descent with negative gradients from incorrect example sampling to refine prompts. It iteratively applies these gradients to address prompt weaknesses, expanding the candidate pool using Monte-Carlo sampling to generate paraphrased versions with synonyms or semantically similar variations. This ensures candidate diversity while guiding the optimization process.
4.1.2 Benchmarks
ProTeGi has been evaluated on benchmark datasets, such as the Liar Dataset Wang (2017) for fake news detection and the Ethos Dataset Mollas et al. (2022) for hate speech detection, which test its ability to handle diverse tasks. Our method will be evaluated on the same benchmarks for comparison. Despite utilizing the publicly available code from ProTeGi to replicate the baseline, we were unable to reproduce the level of metric performance for their method on the Liar Dataset as reported in the original paper.
4.2 Analysis
The initial prompt itself has baseline F1 scores of 0.475 and 0.9 for Liar and Ethos, respectively, evaluated on our test set, serving as the starting point for optimization.
Efficiency. Figures 2, 3, and 4 show the significant efficiency gains of MAPO over ProTeGi. In Figure 2 for Liar, ProTeGi takes approximately 735 seconds to reach an F1 score of 0.58, while MAPO achieves the same score in only 285 seconds. For Ethos, ProTeGi takes 686 seconds while MAPO takes 136 seconds to reach the same level of performance. Overall, there is a 72.74% reduction in runtime. This is notable since ProTeGi’s runtimes can extend into hours (Pryzant et al., 2023), highlighting the resource-intensive nature of automatic prompt optimization.
Figure 3 illustrates that ProTeGi uses about 417 API calls to reach its final F1 score, whereas MAPO requires an average of 77 to reach the same performance, resulting in an average 81.53% reduction. MAPO also completes all six rounds of beam search with 105 fewer API calls on average, addressing critical computational constraints in large-scale optimization.
Figure 4 reinforces these efficiency gains by showing that MAPO surpasses ProTeGi’s performance after just 2 or 3 optimization steps, while ProTeGi requires 6 steps for lower performance. This reduction in steps not only saves time but also suggests a more robust optimization mechanism in MAPO.
While MAPO incurs longer processing times per iteration due to more complex prompt structures and the inclusion of positive gradient history, the overall reduction in steps and runtime demonstrates that these trade-offs do not detract from its overall efficiency.
Efficacy. Figure 4 shows that MAPO consistently outperforms ProTeGi at every beam search round for both Liar and Ethos. MAPO’s F1 score steadily increases while ProTeGi quickly converges and then plateaus or slightly declines. This results in a notable 5.37% increase in overall performance for MAPO, underscoring its superior efficacy in optimizing model outputs.
MAPO’s consistent improvement highlights its effectiveness in leveraging positive gradients and the momentum-based extension to thoroughly explore and refine the prompt search space. By incorporating momentum, MAPO maintains a consistent direction in the optimization process, helping it avoid local minima and erratic updates that hinder progress. This leads to a more robust optimization mechanism compared to ProTeGi. In contrast, ProTeGi’s performance dips suggest it struggles with optimization challenges like local minima and oscillations due to the lack of a stabilizing mechanism. MAPO’s use of momentum and positive gradient history allows it to maintain steady progress, effectively "remembering" beneficial adjustments and reducing the likelihood of stagnation. This results in more reliable convergence toward higher performance levels, demonstrating the superior efficacy of MAPO’s optimization strategy.
Momentum Ablation. In our momentum ablation study, we compared the convergence time required to reach peak ProTeGi performance with and without the use of momentum, as well as the peak MAPO F1 score performance. Table 1 shows that the peak F1 Score of MAPO remains essentially identical regardless of the inclusion of momentum, indicating that momentum does not compromise model accuracy.
However, there is a significant difference in convergence speed: on average, we observe a 54% decrease in convergence time when incorporating momentum into MAPO. This substantial improvement in convergence speed demonstrates that momentum effectively enhances the optimization process by smoothing the loss landscape and helping the model avoid getting trapped in local minima. This is further evidenced by our smoother test set curves depicted in Figures 2, 3, and 4, which contrast sharply with the oscillations observed in ProTeGi’s data. The lack of significant fluctuations in our graphs confirms that our method not only maximizes convergence speed but also promotes stability during training. Collectively, these findings validate the efficiency of incorporating momentum in prompt optimization, proving that our approach accelerates convergence while maintaining, or even slightly improving, the model’s peak performance.
5 Conclusion
In this work, we introduced Momentum-Aided Prompt Optimization (MAPO), a novel momentum-aided extension of natural language gradient descent for prompt optimization in LLMs. Building on ProTeGi, MAPO uses positive natural language gradients and momentum to refine prompts more effectively, guiding optimization consistently and avoiding local minima.
Experiments on the Liar and Ethos datasets show that MAPO outperforms ProTeGi, achieving a 72.74% reduction in convergence time while improving peak F1 scores with fewer API calls and smoother convergence.
Momentum is crucial in overcoming ProTeGi’s limitations like local minima and oscillations. By leveraging gradient history, MAPO ensures a more stable, directed search, resulting in faster convergence and a more robust optimization process.